### v1

@@ Appendix B Omitted Proofs of Section 4
To prove $\operatorname{pqn}(G^{\prime})=1$, we show that given any edge-weight function $w^{\prime}$ for the edges of $G^{\prime}$, there is a PQ-layout of $G^{\prime}$ with one priority queue. From $\langle G^{\prime},w^{\prime}\rangle$, we construct the edge-weighted graph $\langle G,w\rangle$ where $w$ is an edge-weight function for $V$ that is identical [^6] For edges incident to $u$ and $v$, we use the edge weights of the edges incident to $x_{uv}$. So, if there are two edges in $G$ that are represented by the same edge in $G^{\prime}$, they get the same weight. to $w^{\prime}$ except for having a weight of $W$ for edge $e_{0}$ such that $W$ is smaller than any other edge weight in $w$. Since $\operatorname{pqn}(G)=1$, there is a PQ-layout $\Gamma$ with one priority queue for $\langle G,w\rangle$. We next adjust $\Gamma$ slightly to obtain a PQ-layout $\Gamma^{\prime}$ with one priority queue for $\langle G^{\prime},w^{\prime}\rangle$. For an illustration of this process, see Figure 8.
(a) PQ-layout $\Gamma$ before contracting the edge $uv$.
(b) PQ-layout $\Gamma^{\prime}$ after contracting the edge $uv$.
Figure 8: Example of the procedure for maintaining a PQ-layout with one priority queue after edge contraction as described in the proof of Lemma 4.1.
W.l.o.g. let $u$ precede $v$ along the spine of $\Gamma$. Further, let $L$ be the ordered set of vertices to the left of $u$, let $M$ be the ordered set of vertices between $u$ and $v$, and let $R$ be the ordered set of vertices to the right of $v$. In $\Gamma^{\prime}$, we keep $L$ and $R$ unchanged and we place $M\cup\{x_{uv}\}$ in between as follows. First place $x_{uv}$, then add to the right each vertex $m\in M\cap N(x_{uv})$ (where $N(x_{uv})$ are the neighbors of $x_{uv}$) in increasing order with respect to the weight of its incident edge $x_{uv}m$, and finally add the remaining vertices of $M$ in arbitrary order.
It remains to prove that there is no forbidden configuration in $\Gamma^{\prime}$. No pair of edges having both their right endpoints in $L\cup R$ is involved in a forbidden configuration because otherwise, it would have been involved in a forbidden configuration in $\Gamma$ already. The crucial observation is that, in $\Gamma$, no vertex $m\in M$ is the right endpoint of some edge $e$ because the weight $w(e)$ would be greater than $W$, but $e$ would end below $e_{0}$, which has weight $W$. Therefore, the edges incident to vertices in $M$ have their neighbors only in $R\cup\{v\}$. After merging $u$ and $v$ to $x_{uv}$ and placing $x_{uv}$ to the left of all vertices in $M$, the vertices in $M\cap N(x_{uv})$ have exactly one incoming edge from the left, while the other vertices in $M$ have none. By sorting $M\cap N(x_{uv})$ increasingly by the weights of these incoming edges, we avoid a forbidden configuration. Moreover, no edge having its right endpoint in $M\cup\{x_{uv}\}$ is involved in a forbidden configuration with an edge having its right endpoint in $L\cup R$ because this would have been a forbidden configuration in $\Gamma$ already.
See 4.10
@@
For a leaf $\ell$, let $p(\ell)$ denote its neighbor in $C(L)$ and let $e^{\star}$ be an edge of $C(L)$ such that $w(e^{\star})$ is maximum among all edges of $C(L)$. We remove $e^{\star}$ and every leaf $\ell$ with $w({\ell}p(\ell))>w(e^{\star})$ from $L$ obtaining a caterpillar $T$ where both degree-$1$ vertices of the underlying path $P(T)$ are endpoints of $e^{\star}$. We then find a PQ-layout of $T$ using Lemma 4.6 chosing one of the endpoints of $e^{\star}$ as $r$. Note that for each edge $e$ in $T$ we have that $w(e^{\star})\geq w(e)$. While $e^{\star}$ may be co-active with edges in $e$, we have that its right endpoint succeeds any vertex in $T$, i.e., no forbidden configuration is created when we reinsert $e^{\star}$. It remains to reinsert each leaf $\ell$ with $w({\ell}p(\ell))>w(e^{\star})$. Let $S$ denote the set of such leaves and for $\ell\in S$ let $w(\ell)=w({\ell}p(\ell))$. We insert all vertices of $S$ to the right of $r$ sorted in increasing order of weight. It remains to discuss that for $\ell\in S$, edge ${\ell}p(\ell)$ is not involved in forbidden configurations. Let $uv$ with $u\prec v$ be an co-active edge of ${\ell}p(\ell)$. If $v\not\in S$, we have that $v\prec\ell$ and $w(uv)\leq w(e^{\star})<w({\ell}p(\ell))$, i.e., no forbidden configuration occurs. Otherwise, $v$ and $\ell$ are sorted by their weights in $\prec$ which are equal to $w(uv)$ and $w({\ell}p(\ell))$, respectively. Again no forbidden configuration occurs and the statement follows.
Finding and removing a maximum-weight edge of the cycle and all edges with greater weight can be done in $\mathcal{O}(n)$ time. Arranging the resulting graph using Lemma 4.6 as well. For the subsequent reinsertion of edges, we need to sort the leaves by weight. This requires and can be done in $\mathcal{O}(n\log n)$ time, which is also the resulting running time of the algorithm.
See 4.14
Proof B.3.
@@
Assume without loss of generality that $w(ad)\geq\max\{w(ab),w(bc),w(cd)\}$. Consider the edge $e_{c}$ incident to $c$ on $P(C_{c})$. Assume first, that $w(e_{c})\leq w(ad)$. For $\Box$, we use the layout $d\prec c\prec b\prec a$. Note that only the edge $ad$ nests and pseudo-nests edges of $\Box$, however it is the heaviest edge in $\Box$. Further, we use Lemma 4.4 for laying out $C_{a}$ with leftmost vertex $a$. The resulting PQ-layout of $G$ without $C_{c}$ clearly has priority queue number 1. Now, consider the caterpillar $C_{c}^{\prime}$ obtained from $C_{c}$ by removing the leaves of $c$. We lay out $C_{c}^{\prime}$ using Lemma 4.6 and place all vertices of $C_{c}^{\prime}$ except for $c$ before vertex $d$ in $\prec$. Thus, there is one intersection between $C_{c}^{\prime}$ and $\Box$, namely, between $e_{c}$ and $ad$. We have that $c\prec a$ and $w(e_{c})\leq w(ad)$, thus this creates no forbidden configuration.
@@
Next, assume that $w(e_{c})>w(ad)$. For $\Box$, we use the layout $a\prec b\prec c\prec d$. Again, only edge $ad$ nests and pseudo-nests edges of $\Box$, but still it is the heaviest edge in $\Box$. Further, we use Lemma 4.6 for laying out $C_{a}$ with rightmost vertex $a$. Once more, the resulting PQ-layout of $G$ without $C_{c}$ clearly has priority queue number 1. Again, consider the caterpillar $C_{c}^{\prime}$ obtained from $C_{c}$ by removing the leaves of $c$. We lay out $C_{c}^{\prime}$ using Lemma 4.4 and place all vertices of $C_{c}^{\prime}$ except for $c$ after vertex $d$ in $\prec$. Thus, there is one intersection between $C_{c}^{\prime}$ and $\Box$, namely, between $e_{c}$ and $ad$, and one pseudo-nesting, namely between $e_{c}$ and $cd$. However, we have that $d\prec c$ and $w(e_{c})>w(ad)\geq w(cd)$, thus this creates no forbidden configuration.
@@
Finally, in both cases, we only have to reintroduce the leaves $L_{c}$ of vertex $c$ in $C_{c}$ to complete the layout. For $\ell\in L_{c}$, let $w(\ell)=w({\ell}c)$. We iteratively insert the leaves of $c$ with increasing weight. If $w(\ell)\leq w(ad)$, we place $\ell$ directly before $c$ in $\prec$. Then, edge ${\ell}c$ only creates a nesting with edge $ad$, however $ad$ has heavier weight, so no forbidden configuration is introduced. Otherwise, $w(\ell)>w(ad)$ and we find a position for $\ell$ as follows. Let $C_{r}$ be the caterpillar whose non-root vertices are drawn to the right of the vertices of $\Box$, i.e., in the first case $C_{a}$ and in the second case $C_{c}^{\prime}$. Also let $v^{\star}$ be the leftmost vertex of the caterpillar $C_{r}$ with parent $p(v^{\star})$ in $C_{r}$ for which $w(v^{\star}p(v^{\star}))>w(\ell)$. Then, we place $\ell$ directly before $v^{\star}$ in $\prec$. For a contradiction, assume that this creates a forbidden configuration which clearly must involve edge ${\ell}c$ and another edge $uv$ with $u\prec v$ as our procedure ensures that the edges between $c$ and its leaves which pseudo-nest are not creating forbidden configurations. Observe that $uv$ is not an edge of $\Box$ as all edge weights of $\Box$ are smaller than the weight of edge ${\ell}c$ and $\ell$ is placed to the right of all vertices of $\Box$. Thus $u=p(v)$ and if $v\prec\ell$ and $w({\ell}c)<w(uv)$, we have a contradiction to our choice for the position of $\ell$. Thus, we must have $\ell\prec v$ and $w({\ell}c)>w(uv)$. By our choice of the position of $\ell$, we have that $v^{\star}\prec v$. Since also $u\prec\ell$, it follows that $uv$ and $p(v^{\star})v^{\star}$ are co-active, however, this is a contradiction to the fact that $C_{r}$ is drawn with a priority-queue-number-one layout. The proof follows.
See 4.17
@@
Figure 9: Illustration for the proof of Lemma 4.17.
@@
Let $G=K_{2,3}$ and consult Figure 9 for an illustration. Let $\{u_{1},u_{2}\}$ be the vertex partition with vertices of degree 3 and let $\{v_{1},v_{2},v_{3}\}$ be the vertex partition with vertices of degree 2. Let $w$ be any given edge-weight function and denote by $e_{1}$ the heaviest edge of $\langle G,w\rangle$ and by $e_{2}$ the second heaviest edge. We distinguish three main cases and, for each of them, we describe how to construct a PQ-layout $\Gamma$ such that $\operatorname{pqn}(\Gamma)=1$.
  • Case 1: $e_{1}$ and $e_{2}$ are disjoint. Refer to Figure 9(a). We may assume that $e_{1}=u_{1}v_{1}$ and $e_{2}=u_{2}v_{3}$, and define the vertex ordering as $v_{3}\prec u_{1}\prec v_{2}\prec u_{2}\prec v_{1}$. Since $e_{1}$ and $e_{2}$ are the two heaviest edges, there are neither forbidden nestings nor forbidden pseudo-nestings in $\Gamma$. Also, the only two crossing edges are $e_{1}$ and $e_{2}$, and $w(e_{2})\leq w(e_{1})$. Since $e_{2}$ is pulled from the priority queue before $e_{1}$, there is no forbidden crossing configuration.
    • Case 2.1: $e_{3}$ shares a vertex with $e_{1}$. Refer to Figure 9(b). Assume, w.l.o.g., that $e_{1}=u_{1}v_{3}$, $e_{2}=u_{2}v_{3}$, and $e_{3}=u_{1}v_{1}$. The linear ordering of $\Gamma$ is set as $v_{1}\prec u_{2}\prec v_{2}\prec u_{1}\prec v_{3}$. As for Case 1, there are no forbidden (pseudo-)nesting configurations. Also, the only two crossing edges are $e_{2}$ and $e_{3}$, and $w(e_{3})\leq w(e_{2})$. Since $e_{3}$ is pulled from the queue before $e_{2}$, there is no forbidden crossing configuration.
    • Case 2.2: $e_{3}$ shares a vertex with $e_{2}$. Refer to Figure 9(c). Assume, w.l.o.g., that $e_{1}=u_{1}v_{3}$, $e_{2}=u_{2}v_{3}$, and $e_{3}=u_{2}v_{1}$. We define the linear ordering of $\Gamma$ as $v_{1}\prec u_{1}\prec v_{2}\prec u_{2}\prec v_{3}$. The argument is similar to the previous case.
  • Case 3: $e_{1}$ and $e_{2}$ share a degree-3 vertex. Refer to Figure 9(d). Assume, w.l.o.g., that $e_{1}=u_{1}v_{3}$ and $e_{2}=u_{1}v_{1}$. We define the linear ordering of $\Gamma$ as $v_{3}\prec v_{1}\prec u_{2}\prec v_{2}\prec u_{1}$. As in the previous cases, for the choice of $e_{1}$ and $e_{2}$, there are no forbidden nesting or pseudo-nesting configurations. Also, the only two crossing edges are $e_{2}$ and the edge $v_{3}u_{2}$; since $v_{3}u_{2}$ is pulled before $e_{2}$ from the priority queue, and since $w(v_{3}u_{2})\leq w(e_{2})$, there is no forbidden crossing configuration.
See 4.19
@@
It remains to analyze $\mathcal{F}_{4}$$\mathcal{F}_{8}$. For an illustration, refer to Figure 4.
$\mathcal{F}_{4}$, $\mathcal{F}_{5}$
: By Lemma 4.19 and symmetry, we may assume that $a,b\prec c$. Since each edge not in $\triangle abc$ has a weight strictly between $w(ab)=w(ac)$ and $w(bc)$, no such edge can start weakly left of $\min\{a,b\}$ and end strictly right of $\min\{a,b\}$. In particular, it follows that $d\prec a$ and, since also $w(ed)<w(ad)$, that $e\prec d$. In summary, we have $e\prec d\prec a\prec c$ and $b\prec c$.
Now for $\mathcal{F}_{4}$ it follows that the right endpoint of edge $fg$ must be left of $e$. But then edge $ed$ nests below $af$; a contradiction, as $w(ed)>w(af)$.
For $\mathcal{F}_{5}$, first note that if $a\prec b$, then either edge $bf$ ends below $ac$ (if $f\prec c$) or edge $bc$ ends below $bf$ (if $c\prec f$); either is not allowed. So we have $b\prec a$, and hence $f\prec a$, as argued in the beginning. But now either edge $bf$ ends below $ad$ (if $d\prec b$) or edge $ed$ ends below $ab$ (if $b\prec d$); either is not allowed.
$\mathcal{F}_{6}$, $\mathcal{F}_{7}$
: By Lemma 4.19, $b$ or $c$ is the last vertex of the 4-cycle $abcd$ and $a$ or $c$ is the last vertex of $\triangle abc$. Hence, $a,b,d\prec c$. Moreover, $a\prec d$ since otherwise, $ad$ would end below the lighter edge $cd$. We know that $e\prec c$ because otherwise $ac$ would end below the lighter edge $de$ ($ae$, respectively).
For $\mathcal{F}_{6}$, further $e\prec d$ because if $d\prec e\prec c$, then $de$ would end below the lighter edge $cd$. It remains to place $b$ (we already know $b\prec c$). It cannot be $b\prec d$ because then $de$ would end below the lighter edge $bc$. Hence, we have $a\prec d\prec b\prec c$. Now, however, $ab$ ends below the lighter edge $cd$; a contradiction.
For $\mathcal{F}_{7}$, a similar argument applies. We have $e\prec a$ because if $a\prec e\prec c$, then $ae$ would end below one of the lighter edges $ad$ or $cd$. If $b\prec a\prec d\prec c$, then $ae$ would end below the lighter edge $bc$. If $a\prec b\prec d\prec c$, then $ab$ would end below the lighter edge $ad$. Hence, we have $a\prec d\prec b\prec c$. Now, however, $ab$ ends below the lighter edge $cd$; a contradiction.
$\mathcal{F}_{8}$
: By Lemma 4.19, the last vertex of the 4-cycle $abcd$ is $a$ or $b$. In the former case, also $c\prec d$, as otherwise $cd$ ends below the lighter edge $ad$, and $b\prec d$, as otherwise $cd$ ends below the lighter edge $bc$. This leaves four possible orderings: $b\prec c\prec d\prec a$, $c\prec b\prec d\prec a$, and, symmetrically if $b$ is the last vertex, $a\prec d\prec c\prec b$, and $d\prec a\prec c\prec b$.
@@
First suppose that $b\prec c\prec d\prec a$. Then $c\prec f\prec a$ because $df$ has the lowest weight and there are heavier edges ending at $c$ and $a$. Now $f$ cannot be the right endpoint of $fg$ because then the edge $fg$ would end below one of the lighter edges $cd$ or $ad$. However, $f$ can also not be the left endpoint of $fg$ because then the edge $fg$ would end below the lighter edges $cd$ or $ad$, or the heavier edge $ab$ would end below $fg$.

### v3

@@ Appendix B Omitted Proofs of Section˜4
For a leaf $\ell$, let $p(\ell)$ denote its neighbor in $C(L)$ and let $e^{\star}$ be an edge of $C(L)$ such that $w(e^{\star})$ is maximum among all edges of $C(L)$. We remove $e^{\star}$ and every leaf $\ell$ with $w({\ell}p(\ell))>w(e^{\star})$ from $L$ obtaining a caterpillar $T$ where both degree-$1$ vertices of the underlying path $P(T)$ are endpoints of $e^{\star}$. We then find a PQ-layout of $T$ using Lemma˜4.3 chosing one of the endpoints of $e^{\star}$ as $r$. Note that for each edge $e$ in $T$ we have that $w(e^{\star})\geq w(e)$. While $e^{\star}$ may be co-active with edges in $e$, we have that its right endpoint succeeds any vertex in $T$, i.e., no forbidden configuration is created when we reinsert $e^{\star}$. It remains to reinsert each leaf $\ell$ with $w({\ell}p(\ell))>w(e^{\star})$. Let $S$ denote the set of such leaves and for $\ell\in S$ let $w(\ell)=w({\ell}p(\ell))$. We insert all vertices of $S$ to the right of $r$ sorted in increasing order of weight. It remains to discuss that for $\ell\in S$, edge ${\ell}p(\ell)$ is not involved in forbidden configurations. Let $uv$ with $u\prec v$ be an co-active edge of ${\ell}p(\ell)$. If $v\not\in S$, we have that $v\prec\ell$ and $w(uv)\leq w(e^{\star})<w({\ell}p(\ell))$, i.e., no forbidden configuration occurs. Otherwise, $v$ and $\ell$ are sorted by their weights in $\prec$ which are equal to $w(uv)$ and $w({\ell}p(\ell))$, respectively. Again no forbidden configuration occurs and the statement follows.
Finding and removing a maximum-weight edge of the cycle and all edges with greater weight can be done in $\mathcal{O}(n)$ time. Arranging the resulting graph using Lemma˜4.3 as well. For the subsequent reinsertion of edges, we need to sort the leaves by weight. This requires and can be done in $\mathcal{O}(n\log n)$ time, which is also the resulting running time of the algorithm.
See 4.11
@@
First, consider the case that we have a 3-cycle $\triangle=abc$. Suppose that $w(ac)>\max\{w(ab),w(bc)\}$. Then, arrange the vertices of $\triangle$ in order $a\prec b\prec c$. Clearly, this does not lead to any forbidden configuration of $\triangle$. Lay out $C_{a}$ using Lemma˜4.3 such that $a$ is the rightmost vertex of $C_{a}$. Lay out $C_{c}$ using Lemma˜4.1 such that $c$ is the leftmost vertex of $C_{c}$. Combine the PQ-layouts of $\triangle$, $C_{a}$, and $C_{c}$ such that all vertices of $C_{a}$ precede all vertices of $\triangle$ (except for $a$) and that all vertices of $C_{c}$ succeed all vertices of $\triangle$ (except for $c$). Clearly, this is a valid layout with priority queue number 1 because the edge of $C_{c}$ and $\triangle$ are not co-active.
Otherwise, there is an edge in $\triangle$ that is at least as heavy as $ac$. Without loss of generality, let $w(ab)\geq\max\{w(ac),w(bc)\}$. We handle this case as if we had a 4-cycle $\Box$ where we can safely ignore vertex $d$ (having the edge $ac$ instead of a path via the edges $ad$ and $cd$). It is easy to see that the rest of the proof applies to this case.
@@
Second, consider the case that we have a 4-cycle $\Box=abcd$. Assume, without loss of generality, that $w(ab)\geq\max\{w(ad),w(bc),w(cd)\}$. Consider the edge $e_{c}$ incident to $c$ on $P(C_{c})$. Assume first, that $w(e_{c})\leq w(ab)$. For $\Box$, we use the layout $b\prec c\prec d\prec a$. Note that only the edge $ab$ nests and pseudo-nests edges of $\Box$, however it is the heaviest edge in $\Box$. Further, we use Lemma˜4.1 for laying out $C_{a}$ with leftmost vertex $a$. The resulting PQ-layout of $G$ without $C_{c}$ clearly has priority queue number 1. Now, consider the caterpillar $C_{c}^{\prime}$ obtained from $C_{c}$ by removing the leaves of $c$. We lay out $C_{c}^{\prime}$ using Lemma˜4.3 and place all vertices of $C_{c}^{\prime}$ except for $c$ before vertex $b$ in $\prec$. Thus, there is one intersection between $C_{c}^{\prime}$ and $\Box$, namely, between $e_{c}$ and $ab$. We have that $c\prec a$ and $w(e_{c})\leq w(ab)$, thus this creates no forbidden configuration.
@@
Next, assume that $w(e_{c})>w(ab)$. For $\Box$, we use the layout $a\prec d\prec c\prec b$. Again, only edge $ab$ nests and pseudo-nests edges of $\Box$, but still it is the heaviest edge in $\Box$. Further, we use Lemma˜4.3 for laying out $C_{a}$ with rightmost vertex $a$. Once more, the resulting PQ-layout of $G$ without $C_{c}$ clearly has priority queue number 1. Again, consider the caterpillar $C_{c}^{\prime}$ obtained from $C_{c}$ by removing the leaves of $c$. We lay out $C_{c}^{\prime}$ using Lemma˜4.1 and place all vertices of $C_{c}^{\prime}$ except for $c$ after vertex $b$ in $\prec$. Thus, there is one intersection between $C_{c}^{\prime}$ and $\Box$, namely, between $e_{c}$ and $ab$, and one pseudo-nesting, namely between $e_{c}$ and $bc$. However, we have that $w(e_{c})>w(ab)\geq w(bc)$, thus this creates no forbidden configuration.
@@
Finally, in both cases where we consider $\Box$, we only have to reintroduce the leaves $L_{c}$ of vertex $c$ in $C_{c}$ to complete the layout. For $\ell\in L_{c}$, let $w(\ell)=w({\ell}c)$. We iteratively insert the leaves of $c$ with increasing weight. If $w(\ell)\leq w(ab)$, we place $\ell$ directly before $c$ in $\prec$. Then, edge ${\ell}c$ only creates a nesting with edge $ab$, but $ab$ has heavier weight, so no forbidden configuration is introduced. Otherwise, $w(\ell)>w(ab)$ and we find a position for $\ell$ as follows. Let $C_{r}$ be the caterpillar whose non-root vertices are drawn to the right of the vertices of $\Box$, i.e., in the first case $C_{a}$ and in the second case $C_{c}^{\prime}$. Also let $v^{\star}$ be the leftmost vertex of the caterpillar $C_{r}$ with parent $p(v^{\star})$ in $C_{r}$ for which $w(v^{\star}p(v^{\star}))>w(\ell)$. Then, we place $\ell$ directly before $v^{\star}$ in $\prec$. For a contradiction, assume that this creates a forbidden configuration which clearly must involve edge ${\ell}c$ and another edge $uv$ with $u\prec v$ as our procedure ensures that the edges between $c$ and its leaves which pseudo-nest are not creating forbidden configurations. Observe that $uv$ is not an edge of $\Box$ as all edge weights of $\Box$ are smaller than the weight of edge ${\ell}c$ and $\ell$ is placed to the right of all vertices of $\Box$. Thus $u=p(v)$ and if $v\prec\ell$ and $w({\ell}c)<w(uv)$, we have a contradiction to our choice for the position of $\ell$. Thus, we must have $\ell\prec v$ and $w({\ell}c)>w(uv)$. By our choice of the position of $\ell$, we have that $v^{\star}\prec v$. Since also $u\prec\ell$, it follows that $uv$ and $p(v^{\star})v^{\star}$ are co-active, however, this is a contradiction to the fact that $C_{r}$ is drawn with a priority-queue-number-one layout. The proof follows.
See 4.13
@@
Figure 8: Illustration for the proof of Lemma˜4.13. The thickness of an edge indicates the weight of that edge.
Proof B.3.
Let $G=K_{2,3}$ and consult Figure˜8 for an illustration. Let $\{u_{1},u_{2}\}$ be the vertex partition with vertices of degree 3 and let $\{v_{1},v_{2},v_{3}\}$ be the vertex partition with vertices of degree 2. Let $w$ be any given edge-weight function and denote by $e_{1}$ the heaviest edge of $\langle G,w\rangle$ and by $e_{2}$ the second heaviest edge. We distinguish three main cases and, for each of them, we describe how to construct a PQ-layout $\Gamma$ such that $\operatorname{pqn}(\Gamma)=1$.
  • Case 1: $e_{1}$ and $e_{2}$ are disjoint. Refer to Figure˜8(a). We may assume that $e_{1}=u_{1}v_{1}$ and $e_{2}=u_{2}v_{3}$, and define the vertex ordering as $v_{3}\prec u_{1}\prec v_{2}\prec u_{2}\prec v_{1}$. Since $e_{1}$ and $e_{2}$ are the two heaviest edges, there are neither forbidden nestings nor forbidden pseudo-nestings in $\Gamma$. Also, the only two crossing edges are $e_{1}$ and $e_{2}$, and $w(e_{2})\leq w(e_{1})$. Since $e_{2}$ is pulled from the priority queue before $e_{1}$, there is no forbidden crossing configuration.
    • Case 2.1: $e_{3}$ shares a vertex with $e_{1}$. Refer to Figure˜8(b). Assume, w.l.o.g., that $e_{1}=u_{1}v_{3}$, $e_{2}=u_{2}v_{3}$, and $e_{3}=u_{1}v_{1}$. The linear ordering of $\Gamma$ is set as $v_{1}\prec u_{2}\prec v_{2}\prec u_{1}\prec v_{3}$. As for Case 1, there are no forbidden (pseudo-)nesting configurations. Also, the only two crossing edges are $e_{2}$ and $e_{3}$, and $w(e_{3})\leq w(e_{2})$. Since $e_{3}$ is pulled from the queue before $e_{2}$, there is no forbidden crossing configuration.
    • Case 2.2: $e_{3}$ shares a vertex with $e_{2}$. Refer to Figure˜8(c). Assume, w.l.o.g., that $e_{1}=u_{1}v_{3}$, $e_{2}=u_{2}v_{3}$, and $e_{3}=u_{2}v_{1}$. We define the linear ordering of $\Gamma$ as $v_{1}\prec u_{1}\prec v_{2}\prec u_{2}\prec v_{3}$. The argument is similar to the previous case.
  • Case 3: $e_{1}$ and $e_{2}$ share a degree-3 vertex. Refer to Figure˜8(d). Assume, w.l.o.g., that $e_{1}=u_{1}v_{3}$ and $e_{2}=u_{1}v_{1}$. We define the linear ordering of $\Gamma$ as $v_{3}\prec v_{1}\prec u_{2}\prec v_{2}\prec u_{1}$. As in the previous cases, for the choice of $e_{1}$ and $e_{2}$, there are no forbidden nesting or pseudo-nesting configurations. Also, the only two crossing edges are $e_{2}$ and the edge $v_{3}u_{2}$; since $v_{3}u_{2}$ is pulled before $e_{2}$ from the priority queue, and since $w(v_{3}u_{2})\leq w(e_{2})$, there is no forbidden crossing configuration.
See 4.14
(a) Case 1
(b) Case 2.1
(c) Case 2.2
(d) Case 2.3
(e) Case 2.4.1
(f) Case 2.4.2
(g) Case 2.4.3
Figure 9: Illustration for the proof of Lemma˜4.14. The thickness of an edge indicates the weight of that edge.
@@
Let $G$ be $K_{4}$ minus an edge and consult Figure˜9 for an illustration. Let $\{u_{1},u_{2}\}$ be the vertices of degree 3 and let $\{v_{1},v_{2}\}$ be the vertices of degree 2. Let $w$ be any given edge-weight function and denote by $e_{1}$ the heaviest edge of $\langle G,w\rangle$, by $e_{2}$ the second heaviest edge, and by $e_{3}$ the third heaviest edge. We distinguish two main cases and, for the second case, four subcases. We describe how to construct a PQ-layout $\Gamma$ such that $\operatorname{pqn}(\Gamma)=1$.
  • Case 1: $e_{1}=u_{1}u_{2}$. Refer to Figure˜9(a). Assume, without loss of generality (due to symmetry), that $e_{2}=u_{1}v_{1}$. We define the vertex ordering as $v_{1}\prec u_{2}\prec v_{2}\prec u_{1}$. The only edges spanning over other edges are $e_{1}$ and $e_{2}$. Since they are the heaviest edges and end at the rightmost vertex, there is no forbidden configuration with another edge. Furthermore, $e_{1}$ and $e_{2}$ together do not create a forbidden configuration since they share a common right endpoint.
  • Case 2: $e_{1}=u_{1}v_{1}$. Note that this is the only remaining case to assign $e_{1}$ since all edges except for $u_{1}u_{2}$ are equivalent by the symmetry of $G$. We consider four subcases for the four choices of the edge $e_{2}$, and, in the last subcase, we distinguish further between the choices for edge $e_{3}$.
    • Case 2.1: $e_{2}=u_{1}u_{2}$. Refer to Figure˜9(b). As in Case 1, set the linear ordering of the vertices to $v_{1}\prec u_{2}\prec v_{2}\prec u_{1}$. Again, the only edges spanning over other edges are $e_{1}$ and $e_{2}$, and $e_{1}$ and $e_{2}$ share a common right endpoint.
    • Case 2.2: $e_{2}=u_{2}v_{2}$. Refer to Figure˜9(c). Set the linear ordering of the vertices to $v_{2}\prec u_{1}\prec u_{2}\prec v_{1}$. Again, the only edges spanning over other edges are $e_{1}$ and $e_{2}$. Now, however, $e_{1}$ and $e_{2}$ intersect, but they do not create a forbidden configuration because $w(e_{1})>w(e_{2})$ and $e_{1}$ ends after $e_{2}$.
    • Case 2.3: $e_{2}=u_{1}v_{2}$. Refer to Figure˜9(d). Set the linear ordering of the vertices to $v_{1}\prec v_{2}\prec u_{2}\prec u_{1}$. Since the edges $e_{1}$ and $e_{2}$ are the heaviest edges, end at the rightmost vertex, and share a common right endpoint, they are not involved in any forbidden configuration. The only remaining pair of co-active edges is $u_{2}v_{1}$ and $u_{2}v_{2}$. Since they share $u_{2}$ as their common right endpoint, they do not induce a forbidden configuration.
    • Case 2.4.1: $e_{2}=u_{2}v_{1}$ and $e_{3}=u_{1}u_{2}$. Refer to Figure˜9(e). Set the linear ordering of the vertices to $u_{1}\prec v_{2}\prec u_{2}\prec v_{1}$. Note that $e_{1}$ spans over all edges but also has the largest weight, so $e_{1}$ is not involved in a forbidden configuration. The only remaining edge that spans over other edges is $e_{3}$. However, note that both $u_{1}v_{2}$ and $u_{2}v_{2}$ do not have a greater weight than $e_{3}$, so there is also no forbidden configuration.
    • Case 2.4.2: $e_{2}=u_{2}v_{1}$ and $e_{3}=u_{1}v_{2}$. Refer to Figure˜9(f). Set the linear ordering of the vertices to $u_{1}\prec u_{2}\prec v_{2}\prec v_{1}$. The two heaviest edges ($e_{1}$ and $e_{2}$) have the rightmost endpoint, so they are not involved in a forbidden configuration. No edge except for $e_{1}$ and $e_{2}$ end to the right of the third heaviest edge $e_{3}$, so $e_{3}$ is not involved in a forbidden configuration. The other two remaining edges are not co-active.
    • Case 2.4.3: $e_{2}=u_{2}v_{1}$ and $e_{3}=u_{2}v_{2}$. Refer to Figure˜9(g). Set the linear ordering of the vertices to $u_{2}\prec u_{1}\prec v_{2}\prec v_{1}$. There are no forbidden configurations by the same argument as in Case 2.4.2; the only difference is that $e_{1}$ and $e_{2}$ are swapped, but both end at the same right endpoint.
See 4.15
@@
It remains to analyze $\mathcal{F}_{4}$$\mathcal{F}_{8}$. For an illustration, refer to Figure˜10.
$\mathcal{F}_{4}$
: We consider the cases that $G=\mathcal{F}_{4}$ and that $G$ is larger graph that has $\mathcal{F}_{4}$ as a minor both at once. With a sequence of vertex splits (potentially after renaming), we can (i) replace the edge $bc$ by a path (we denote a sequence of additional vertices by $x^{*}$), or (ii) replace the edge $ad$ by a path (we denote a sequence of additional vertices by $y^{*}$). In any case, we assign any additional edge a weight of 2.
(a) $\mathcal{F}_{1}$
(b) $\mathcal{F}_{2}$
(c) $\mathcal{F}_{3}$
(d) $\mathcal{F}_{4}$
(e) $\mathcal{F}_{5}$
(f) $\mathcal{F}_{6}$
(g) $\mathcal{F}_{7}$
(h) $\mathcal{F}_{8}$
Figure 10: (Copy of Figure˜4.) Forbidden minors for graphs with priority queue number 1. The graphs with the provided edge weights do not admit a linear layout with priority queue number 1. The thickness of an edge indicates the weight of that edge.
By Lemma˜4.15, the last vertex of the cycle $abx^{*}cdy^{*}$ is $a$ or $b$. If $a$ is the last vertex, then also $c\prec d$, as otherwise $cd$ ends below a lighter edge of weight 2 on the path from $d$ to $a$. Furthermore, $b\prec d$, as otherwise $cd$ ends below a lighter edge of weight 2 on the path between $b$ and $d$. The symmetric arguments apply when we swap the roles of $a$ and $b$ and simultaneously the roles of $c$ and $d$.
This leaves four possible orderings: If $a$ is the last vertex, we have $b\prec c\prec d\prec a$ or $c\prec b\prec d\prec a$. If $b$ is the last vertex, we have $a\prec d\prec c\prec b$ or $d\prec a\prec c\prec b$.
@@
First, suppose that $b\prec c\prec d\prec a$. No vertex from $x^{*}$ on the path between $b$ and $c$ lies to the right of $d$ since then an edge of weight 2 spans over the end of the heavier edge $cd$ of weight 3. Observe that $c\prec f\prec a$ because $df$ has the lowest weight and there are heavier edges ending at $a$ and at $c$ and $x^{*}$. Now $f$ cannot be the right endpoint of $fg$ because then the edge $fg$ would end below the lighter edge $cd$ or below a lighter edge on the path from $d$ to $a$. However, $f$ can also not be the left endpoint of $fg$ because then, if $f\prec g\prec a$, the edge $fg$ would end below the lighter edge $cd$ or below a lighter edge on the path from $d$ to $a$, or, if $f\prec a\prec g$, the heavier edge $ab$ would end below $fg$.